"Find the equation of a line that passes through (3, -5) and whose slope is -(3/7)". I really need this ASAP. Thanks.
Answer by ewatrrr(10682)

 
Hi
line that passing through (3, -5) with slope = -(3/7)
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
y = (-3/7)x + b
-5 = -3/7*3 + b |using ordered pair (3,-5) to solve for b
-35/7 +9/7 = b
-26/7 = b
y = (-3/7)x -26/7 OR 3x+7y = -26 in standard form